How do u turn x-2Y=3 into slope intercept form
Answer by stanbon(75887)

turn x-2Y=3 into slope intercept form
-----
Solve for "y"::
2y = x-3
-----
y = (1/2)x - (3/2)
-----
slope = 1/2
intercept = -3/2
------------
Cheers,
Stan H.